### v1

# Manturov Projection for Virtual Legendrian Knots in $ST^{*}F$

Abstract.
Kauffman virtual knots are knots in thickened surfaces $F\times\mathbb{R}$ considered up to isotopy, stabilizations and destabilizations, and diffeomorphisms of $F\times\mathbb{R}$ induced by orientation preserving diffeomorphisms of $F$. Similarly, virtual Legendrian knots, introduced by Cahn and Levi [4], are Legendrian knots in $ST^{*}F$ with the natural contact structure. Virtual Legendrian knots are considered up to isotopy, stabilization and destabilization of the surface away from the front projection of the Legendrian knot, as well as up to contact isomorphisms of $ST^{*}F$ induced by orientation preserving diffeomorphisms of $F$.
We show that there is a projection operation $\mathop{\mathrm{proj}}\nolimits$ from the set of virtual isotopy classes of Legendrian knots to the set of isotopy classes of Legendrian knots in $ST^{*}S^{2}$. This projection is obtained by substituting some of the classical crossings of the front diagram for a virtual crossing. It restricts to the identity map on the set of virtual isotopy classes of classical Legendrian knots. In particular, the projection $\mathop{\mathrm{proj}}\nolimits$ extends invariants of Legendrian knots to invariants of virtual Legendrian knots. Using the projection $\mathop{\mathrm{proj}}\nolimits$, we show that the virtual crossing number of every classical Legendrian knot equals its crossing number. We also prove that the virtual canonical genus of a Legendrian knot is equal to the canonical genus.
The construction of $\mathop{\mathrm{proj}}\nolimits$ is inspired by the work of Manturov.
1. Introduction
A Gauss Diagram of a knot $K$ in $\mathbb{R}^{3}$ is a circle parametrizing the knot. For each double point in the knot diagram, the two preimages on the circle are connected by a chord oriented from the preimage of the upper strand to the preimage of the lower strand. These chords are equipped with the sign $\pm$ corresponding to the sign of the double point in the diagram. The Reidemeister moves for knot diagrams can be expressed in terms of Gauss Diagram moves, as described by Polyak and Viro in [24]. The equivalence classes of Gauss Diagrams modulo these moves are virtual knots, see Kauffman [15].
As proved by Goussarov, Polyak, and Viro [13] (see also Kauffman [14]), two different classes of knots do not become equivalent in the virtual category. Carter, Kamada, Saito [5] proved that virtual knots can be viewed as equivalence classes of knots in a thickened surface $F\times\mathbb{R}$ modulo isotopies and stabilizations/destabilizations of $F$ (adding and deleting handles) away from the knot projection, as well as automorphisms of $F\times\mathbb{R}$ of the form $\alpha\times\mathop{\mathrm{id}}\nolimits_{\mathbb{R}}$ where $\alpha$ is an orientation preserving diffeomorphism of $F$.
We say that a knot $K$ in a thickened surface $F\times K$ is irreducible if there is no destabilization of $F$ away from the projection of $K$ to $F$. Given a knot $K$ in a thickened surface $F\times\mathbb{R}$, the Kuperberg theorem [16] asserts that $K$ is virtually isotopic to an irreducible knot $K^{\prime}$ in a thickened surface $F^{\prime}\times\mathbb{R}$. Furthermore, the pair $(F^{\prime}\times\mathbb{R},K^{\prime})$ is unique up to an automorphism of $F^{\prime}\times\mathbb{R}$ induced by orientation preserving automorphism of $F^{\prime}$, and the surface $F^{\prime}$ can be obtained from $F$ using only destabilizations and orientation preserving automorphisms.
A contact form on an odd dimensional manifold $Z^{2k+1}$ is a $1$-form $\alpha$ such that $\alpha\wedge d\alpha^{k}$ is nowhere zero. A cooriented contact structure on $Z^{2k+1}$ is a cooriented (transversally oriented) hyperplane distribution such that it is the kernel of some contact form with the coorientation pointing into the half spaces where $\alpha$ is positive.
A contact structure is characterized as maximally non-integrable hyperplane distribution, meaning that the maximum dimension of an everywhere tangent submanifold is $k$ in $(Z^{2k+1},\ker\alpha)$. Such a submanifold $L^{k}$ is called Legendrian.
The spherical cotangent bundle $ST^{*}M$ of $M$ is the space of all nonzero linear functionals $T_{p}M\to\mathbb{R}$ considered up to multiplication by a positive scalar. It is equipped with the natural projection $\pi:ST^{*}M\to M$. The space $ST^{*}M$ has a natural contact structure where the contact plane at a linear functional $l\in ST^{*}M$ is $d\pi^{-1}(\ker l)$.
A Legendrian submanifold $L\subset ST^{*}M$ is fully described by its naturally cooriented front projection $\pi(L)\subset M$. Note that in general the front of $L$ is not immersed and has various Legendrian singularities such as swallow tails etc.
Virtual Legendrian knots were introduced by Cahn and Levi [4]. These are Legendrian knots $L$ in $ST^{*}F$, where $F$ is a surface (or more generally in $ST^{*}M$, where $M$ is a higher dimensional manifold), considered modulo Legendrian isotopies and modifications of $ST^{*}F$ induced by stabilization and destabilization of $F$ away from the front projection of $L$, as well as contact automorphisms of $ST^{*}F$ induced by orientation preserving diffeomorphisms of $F$. They showed that all the alternative view points of the classical virtual knot theory in $F\times\mathbb{R}$ also hold for virtual Legendrian knots in $ST^{*}F$. That is they can be viewed as equivalence classes of formal generalized Gauss diagrams with special symbols for various cusps and crossing points of the front projection and they can be viewed as cooriented front diagrams with extra virtual crossings.
The authors of this paper proved a generalization of the Kuperberg theorem to virtual Legendrian knots [11] and, in particular, concluded that two non-isotopic Legendrian knots in $ST^{*}S^{2}$ are also not virtually Legendrian isotopic.
We also applied [12] the virtual Legendrian knot theory to the study of causality in Borde-Sorkin spacetimes [3, 25] and made a few related conjectures in the spirit of the works of the first author and Nemrovski [7, 8, 9, 10] that were motivated by the Low Conjecture [17, 18, 19, 20] and the Legendrian Low Conjecture of Natario and Tod [22] from the Penrose school of General Relativity [23].
2. Projection and Parities for Virtual Knots in $ST^{*}F$
This section is largely motivated by the ideas of the Manturov projection [21] of virtual knots in $F\times\mathbb{R}.$
Consider an irreducible Legendrian knot $L$ in $ST^{*}F$. Let $D(L)$ be the diagram of $L$. We will denote the classical intersection points of $D(L)$ by the same letters as the intersection points of the front diagram $\pi(L)$ of $L$. Each classical self-crossing $p$ of $\pi(L)$ separates $\pi(L)$ into two loops that are obtained from $\pi(L)$ by choosing an orientation of $L$, and then smoothing $\pi(L)$ at $p$ according to the orientation of $L$. We call the self-crossing $p$ of $\pi(L)$ odd if each crossing $p$ of $\pi(L)$ separates it into two loops and if it is not true that the classes of both loops in $\pi_{1}(F,p)$ are powers of the same third element, then $p$ is odd. Note that these powers can be positive, zero or negative.
Manturov projection of a virtual irreducible Legendrian knot $L$ in $ST^{*}F$ is the classical Legendrian knot $\mathop{\mathrm{proj}}\nolimits(L)$ such that the diagram $D(\mathop{\mathrm{proj}}\nolimits(L))$ is obtained from $D(L)$ by changing each odd crossing of $\pi(L)$ to a virtual crossing and keeping even crossings of $\pi(L)$ intact. If $L$ is a virtual Legendrian knot that is not necessarily irreducible, then, by our generalization [11, Theorem 1.1] of the Kuperberg [16] Theorem, virtual isotopy class of $L$ has a unique irreducible representative $L^{\prime}$, defined up to Legendrian isotopy and automorphism of $F$. We define $\mathop{\mathrm{proj}}\nolimits(L)$ to be $\mathop{\mathrm{proj}}\nolimits(L^{\prime})$.
Theorem 2.1.
The Legendrian isotopy class of the Manturov projection of a virtual Legenrian knot in $ST^{*}F$ does not depend on the representative of a virtual knot and maps virtually Legendrian isotopic knots to isotopic Legendrian knots. In particular, if a virtual Legendrian knot $L_{v}$ is virtual isotopic to a classical Legendrian knot $L$ in $ST^{*}S^{2}$, then its projection $\mathop{\mathrm{proj}}\nolimits(L_{v})$ is Legendrian isotopic to that knot $L$.
The proof of Theorem 2.1 follows immediately from Theorem 3.1.
3. Nikonov theorem
Theorem 3.1.
There is a map $pr$ from diagrams of irreducible virtual Legendrian knots to classical knot diagrams such that for every knot $K$ we have $\mathop{\mathrm{proj}}\nolimits(K)$ is obtained from $K$ by substituting odd crossings to virtual crossings; and if two diagrams $K_{1}$ and $K_{2}$ are related by a Reidemeister move of Legendrian knots (performed within the given minimal genus diagram), then their images $\mathop{\mathrm{proj}}\nolimits(K_{1})$ and $\mathop{\mathrm{proj}}\nolimits(K_{2})$ are related by a Reidemeister move of Legendrian knots.
Proof.
Let $K$ be an irreducible Legendrian knot in $ST^{*}F$ for a surface $F$ of genus $g$. Recall that we denote by $\pi$ the projection of the circle bundle $ST^{*}F$ onto the base surface $F$. If the front $\pi(K)$ of $K$ is generic, then it only has transverse double points and semi-cubical cusps. By identifying $ST^{*}F$ with the set of cooriented contact elements at points of $F$, we deduce a coorientation of the front $\pi(K)$. Let $D(K)$ denote the cooriented diagram in $\mathbb{R}^{2}$ of the virtual Legendrian knot $K$. The diagram $D(K)$ can be obtained from $K$ by the composition of the projection of $K$ to $F\setminus\{pt\}$ for some point $pt$ outside of $\pi(K)$, and an immersion $F\setminus\{pt\}\to\mathbb{R}^{2}$. All crossings of $D(K)$ that do not correspond to crossings of $\pi(K)$ are marked as virtual. We refer to non-virtual crossings as to classical.
Every classical crossing $v$ corresponds to two loops $h_{v,1}$ and $h_{v_{2}}$ of $\pi(K)$ on $F$. We define $\mathop{\mathrm{proj}}\nolimits$ by replacing some of the classical crossings of $D(K)$ with virtual ones. Namely, if the homotopy classes of $h_{v,1}$ and $h_{v,2}$ are powers of the same element of the fundamental group, then the crossing remain classical, and otherwise it becomes virtual. Here we permit the power to be positive, negative, or zero. As in the case of virtual knots, $\mathop{\mathrm{proj}}\nolimits(K)$ is a diagram of a Legendrian knot over the sphere $S^{2}$. Thus, indeed there is a map $\mathop{\mathrm{proj}}\nolimits$.
Suppose that the knot $K^{\prime}$ is obtained from the knot $K$ by a cusp birth move, which corresponds to the first Reidemeister move of knots. Then at least one of the two loops, $h_{v,1}$ or $h_{v,2}$, is trivial. Therefore, the crossing remains classical, which means that $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$ is obtained from $\mathop{\mathrm{proj}}\nolimits(K)$ by the same move.
The second Reidemeister move of knots corresponds to the fold-cusp move (when a cusp pinches a branch) and to the two possible safe self-tangencies when the coorientations of the tangent branches point in opposite directions and hence do not give a double point of $K.$ As in the argument for knots, either the two new crossings in $K^{\prime}$ remain in $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$, or both new crossings in $K^{\prime}$ become virtual in $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$.
The third Reidemeister move of Legendrian knots is similar to the third Reidemeister move for knots. By an argument in the proof of [21, Theorem 6], out of three crossings in $K$ either only one remains classical, or all three remain classical in $\mathop{\mathrm{proj}}\nolimits(K)$. (So the case where exactly one of the crossings becomes virtual is not possible.) If all three crossings remain classical, then $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$ is obtained from $\mathop{\mathrm{proj}}\nolimits(K)$ by the same third Reidemeister move.
Suppose now that only one of the three crossings in $K$ remain classical in $\mathop{\mathrm{proj}}\nolimits(K)$. Then in $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$ also only one of the crossings in classical. Thus, again $\mathop{\mathrm{proj}}\nolimits(K^{\prime})$ is obtained from $\mathop{\mathrm{proj}}\nolimits(K)$ by the third Reidemeister move for virtual knots. ∎
4. Crossing number and canonical genus corollaries of the Projection Theorem
Definition 4.1.
Given a knot diagram $D$ the knot $K$ its canonical genus $g(D)$ is the genus of a surface obtained as follows: smoothen the classical crossings of $D$ using orientation of the knot branches and glue the resulting ovals with disks, then glue the bands to the disks at each crossing. The canonical genus $\mathop{\mathrm{cg}}\nolimits(K)$ is the minimum of $g(D)$ over all the knot diagrams of $D.$ Clearly this procedure gives also the virtual canonical genus $\mathop{\mathrm{vcg}}\nolimits(K_{v})$ of a virtual knot $K_{v}$ with the minimum taken over the class of virtual knot diagrams of $K_{v}.$ It is clear that when the virtual knot $K_{v}$ is isotopic to a classical knot $K$ we have $\mathop{\mathrm{vcg}}\nolimits(K_{v})\leq\mathop{\mathrm{cg}}\nolimits(K)$, see Stoimenov, Tchernov, Vdovina [26]. In [6] the first author used the Manturov Projection Theorem to prove that for a classical knot $K$ we have $\mathop{\mathrm{vcg}}\nolimits(K)=\mathop{\mathrm{cg}}\nolimits(K)$.
It is easy to see that the quantities $g(D)$ and hence $\mathop{\mathrm{vcg}}\nolimits(L_{v})$ and $\mathop{\mathrm{cg}}\nolimits(L)$ are well defined for the classical and virtual Legendrian front diagram and we clearly have the inequality $\mathop{\mathrm{vcg}}\nolimits(L_{v})\leq\mathop{\mathrm{cg}}\nolimits(L).$ Note that the knot $L$ does not have to be zero homologous and hence the actual Seifert surface given by the above procedure does not have to exist. Repeating the argument of [6] but using Projection Theorem 2.1 instead of the Manturov Projection Theorem, we get the following Theorem.
Theorem 4.2.
If a virtual Legendrian knot $L_{v}$ is virtually isotopic to the classical Legendrian knot $L\subset ST^{*}S^{2}$ then $\mathop{\mathrm{vcg}}\nolimits(L_{v})=\mathop{\mathrm{cg}}\nolimits(L).$
Define the crossing number $\mathop{\mathrm{cr}}\nolimits(L)$ of a Legendrian knot in $ST^{*}F$ to be the minimal number of crossing points in its front diagram. Similarly define the virtual crossing number $\mathop{\mathrm{cr}}\nolimits_{v}(L_{v})$ of a virtual Legendrian knot $L_{v}$ to be the minimal number of classical crossing points in its virtual front diagram.
Similar to Belousov [1] definition for classical knots, define the $k$-arc crossing number $\mathop{\mathrm{cr}}\nolimits^{k}(L)$ to be the minimal number of classical crossing numbers among all presentation of $L$ by all diagrams that can be decomposed into $k$ arcs without classical self-intersection points. Similarly, define the virtual $k$-arc crossing number $\mathop{\mathrm{cr}}\nolimits_{v}^{k}(L)$ to be the minimal number of classical crossings among all diagrams of a virtual Legendrian knot $L$ that can be decomposed into $k$ arcs without self-intersection points.
Clearly for a Legendrian knot $L$ we have $\mathop{\mathrm{cr}}\nolimits_{v}(L)\leq\mathop{\mathrm{cr}}\nolimits(L)$ and for all positive $k\in\mathbb{N}$ we have $\mathop{\mathrm{cr}}\nolimits_{v}^{k}(L)\leq\mathop{\mathrm{cr}}\nolimits_{v}(L)$, and $\mathop{\mathrm{cr}}\nolimits^{k+1}_{v}(L)\leq\mathop{\mathrm{cr}}\nolimits^{k}_{v}(L)$ for all positive $k\in\mathbb{N}.$
Theorem 4.3.
For a classical Legendrian knot $L\subset ST^{*}S^{2}$ we have $\mathop{\mathrm{cr}}\nolimits_{v}(L)=\mathop{\mathrm{cr}}\nolimits(L).$ Similarly for a classical Legendrian knot $L$ we have $\mathop{\mathrm{cr}}\nolimits^{k+1}(L)\leq\mathop{\mathrm{cr}}\nolimits^{k}(L)$ and for a Legendrian knot $L$ and positive $k\in\mathbb{N}$ we have $\mathop{\mathrm{cr}}\nolimits^{k}_{v}(L)=\mathop{\mathrm{cr}}\nolimits^{k}(L).$
Proof.
Take a virtual diagram of $L$ with the minimal number of classical crossings. By the projection theorem 2.1 we can erase some of these crossings and get the same classical Legendrian knot, thus we have the other inequality $\mathop{\mathrm{cr}}\nolimits_{v}(L)\geq\mathop{\mathrm{cr}}\nolimits(L).$
The remaining inequality for the $k$-arc virtual and $k$-arc classical crossing number is proved similarly. ∎
Remark 4.4.
For classical knots in $\mathbb{R}^{3}$ the $k$-arc crossing number was introduced by Belousov [1]. The statements similar to Corollary 4.3 for Kauffman rather than Legendrian virtual knots was first proved by Belousov, Chernov, Malyutin and Sadykov [2].